Grade 12Algebra

Find the product, using suitable properties:
8 × 53 × (–125)

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

The given equation is in the form of Commutative law of Multiplication.
=a × b=b ×a
Then,
= 8 × [53 × (-125)]
= 8 × [(-125) × 53]
= [8 × (-125)] × 53
= [-1000] × 53
= - 53000

Thanks